### v1

# The Rise of Small Language Models in Healthcare: A Comprehensive Survey

Abstract.
Despite substantial progress in healthcare applications driven by large language models (LLMs), growing concerns around data privacy, and limited resources; the small language models (SLMs) offer a scalable and clinically viable solution for efficient performance in resource-constrained environments for next-generation healthcare informatics. Our comprehensive survey presents a taxonomic framework to identify and categorize them for healthcare professionals and informaticians. The timeline of healthcare SLM contributions establishes a foundational framework for analyzing models across three dimensions: NLP tasks, stakeholder roles, and the continuum of care. We present a taxonomic framework to identify the architectural foundations for building models from scratch; adapting SLMs to clinical precision through prompting, instruction fine-tuning, and reasoning; and accessibility and sustainability through compression techniques. Our primary objective is to offer a comprehensive survey for healthcare professionals, introducing recent innovations in model optimization and equipping them with curated resources to support future research and development in the field. Aiming to showcase the groundbreaking advancements in SLMs for healthcare, we present a comprehensive compilation of experimental results across widely studied NLP tasks in healthcare to highlight the transformative potential of SLMs in healthcare. The updated repository is available at Github [^1] https://github.com/drmuskangarg/SLMs-in-healthcare/.
carbon emission reduction, mental health analysis, small language models, healthcare informatics
copyright: nonepublicationmonth: 8
1. Background
The global small language model (SLM) market is projected to reach USD 29.64 billion by 2032, with the healthcare industry anticipated to experience the fastest compound annual growth rate (CAGR) of 18.31% from 2024 to 2032222https://www.globenewswire.com/. This growth is driven by the growing adoption of SLM in medical diagnosis, patient care, and administrative processes. The use of SLM in analyzing extensive medical literature and patient data provides clinicians with evidence-based interpretations to support clinical decision making [^3] https://www.healthcare.digital/.
A rapid evolution of transformer-based architectures has catalyzed a paradigm shift in natural language processing (NLP), with large (or very large) language models (LLMs) exceeding hundreds of billions of parameters (e.g., Med-PaLM (Tu et al., 2024)), achieving state-of-the-art performance on various benchmarks (Singhal et al., 2025). However, the computational and energy demands associated with the training, fine-tuning and deployment of LLMs in production require GPU clusters and significant power consumption at the megawatt scale, driving substantial interest in SLMs in both research and healthcare applications (Popov et al., 2025). Healthcare is a broad, overarching system that encompasses the prevention, diagnosis, treatment, and management of illness, along with the preservation of physical and mental well-being through services offered by medical, nursing, and allied health professionals.
1.1. Motivation
The deployment of language models in clinical settings demands an uncertain trade-off among high-stakes accuracy, regulatory compliance, and computational efficiency (Singh et al., 2023). While LLMs effectively capture clinical concepts (Fu et al., 2023), their high inference latency (>500ms per query) and memory demands (>16GB VRAM) make them unsuitable for real-time analysis of medical records with protected health information (PHI). Healthcare data is governed by stringent regulations such as HIPAA and GDPR and requires on-premise processing to mitigate privacy risks (Jonnagaddala and Wong, 2025). Regulatory agencies (e.g., FDA [^4] https://www.fda.gov/) encourages transparent AI-driven decision-making to promote interpretation. Knowledge-driven models generate interpretable results based on trusted external medical knowledge bases, supporting regulatory compliance such as EU Medical Device Regulation (MDR) (Gilbert, 2023). Training LLMs cause an immense carbon footprint of more than 500 metric tons $\text{CO}_{2}$ which is not sustainable for environmental goals. In contrast, SLM is optimized for watts per inference, a sustainability metric (Potkins et al., 2024). They reduce energy consumption by 10 to 100 times through optimization and compression techniques, addressing the Jevons paradox in green AI (making models more energy-efficient can sometimes lead to an overall increase in resource consumption—not a decrease, as one might expect) (Luccioni et al., 2025). Critical care scenarios require subsecond decision making driven by clinical NLP (cNLP) tasks where LLMs with high inference latencies are not suitable for time-sensitive applications (Bai et al., 2024). As such, we opt to expose the potential of SLM to address specialized healthcare challenges under limited resources and time constraints.
1.2. Language model fundamentals
SLM models lie on the "knee" of Pareto frontier, a curve that represents the set of solutions to multi-objective optimization where no objective can be improved without worsening at least one other objective (Cho et al., 2025), maximizing the utility per parameter. These models with fewer parameters than LLMs, aim for Pareto efficiency, a change that makes at least one person better off without making anyone else worse off. This balances performance with minimal use of resources, reducing latency, memory consumption and environmental impact. Consequently, SLMs are optimized for production-ready deployment, which is achieved through architectural innovations and task-aware specializations (Van Nguyen et al., 2024). Unlike LLMs, SLMs in healthcare are designed as democratization (Accessible, participatory, and equitable model development) tools (Subramonian et al., 2024), enabling decision-making in low-resource settings and shaping a design philosophy that extends beyond model size. Platforms such as Pytorch [^5] https://pytorch.org/, Hugging Face [^6] https://huggingface.co/ and Lamini [^7] https://www.lamini.ai/ define the practical deployability of language models that run on consumer hardware without specialized infrastructure.
Transformers overview: SLMs are transformer-based models that capture contextual relationships and dependencies in language. A transformer is a deep learning model that processes sequences in parallel by converting tokenized words into numerical embeddings. Transformers use positional encoding to maintain word order and allow parallel processing. The transformer encoder converts input sequences into context-rich representations, while the decoder generates output sequences. Transformers are categorized into encoder-only, decoder-only, and encoder-decoder architectures. The encoder captures relationships among words using self-attention mechanisms in the attention layer, assigning weights based on contextual relevance (Belhaouari and Kraidia, 2025; Liu and Xu, 2023). Encoder-only models (BERT-like) are generally good in discriminative tasks. Decoder-only models (GPT-like) generate text by predicting the next token in a sequence, making them ideal for generative tasks. Encoder-decoder models (e.g., T5, BART (Lewis et al., 2019)) combine input understanding and output generation. Given our focus on generative tasks in clinical settings, we emphasize decoder-only SLMs for their efficiency and adaptability.
1.3. Scope
Early studies define SLM as the models with 1B to 20B parameters, while LLMs have hundreds of billion parameters (Subramanian et al., 2025). Industry research defines SLMs based on deployment limits, such as models running on CPU-only edge devices (Sun et al., 2020). Later research suggests a relative definition, where SLMs are 10 to 100 times smaller than the largest models of their time (Jovanovic and Campbell, 2024). The definition of “small" in SLMs is relative in time and evolves with hardware advancements. It generally refers to reduced computational complexity (e.g., parameter count) and lower storage requirements (e.g., model size).
Parameter range: SLMs typically fall within a spectrum in which the lower bound represents the minimum size required to exhibit emergent abilities for healthcare applications, while the upper bound is constrained by the maximum size manageable under limited resource conditions (Lu et al., 2024a). While models exceeding 7B parameters typically require high-end GPUs, 7B represents a practical upper limit for local execution; our study defines SLM to be up to 7B parameter range, though future models may surpass this.
Search strategy: The existing literature used mixed terminology for referring to the SLM: local language models, remote language models, compressed models, optimized models, on-device language models, small LLMs and lightweight language models. We used these keywords to identify SLMs designed for NLP-driven clinical decision making from well-known sources: PubMed, ACM, IEEE, Google Scholar, Scopus. Our study covers three dimensions: (i) NLP tasks, (ii) stakeholders, (iii) care continuum. The NLP tasks covers question answering (QA), named entity recognition (NER), relation extraction (RE), information extraction (IE), natural language inference/ understanding (NLI/ NLU), text classification (CLS), summarization (SUM). We focus on three important stakeholders: patient, healthcare professionals, research/material. The care continuum covers triaging, diagnosis, treatment, recovery, risk prediction, and virtual health assistants.
Inclusion criteria. Unlike fundamental bidirectional BERT-like transformers, our interest lies in the decoder-only transformer architecture models (traced back to GPT) in the healthcare domain with open weights. We include SLMs trained on English language data for clinical NLP tasks, developed before April 2025. The scope of our study is defined in Fig. 1(a). Exclusion criteria. We exclude models that are non-English, designed for inferring non-textual audiovisual data, multimodal SLMs, agentic AI systems, and models exceeding 7 billion parameters. Additionally, models lacking peer-reviewed publications for SLM models are not considered in this study.
(a) Scope of our study
(b) Monthly downloads
Figure 1. Opening remarks of the study rationale: (a) Defined scope of the study to investigate the NLP-centered SLM development for healthcare professionals; (b) Monthly downloads for 7B and 13B parameters model from Huggingface for both MentaLLAMA and MedAlpaca.
{forest}
[…]
Figure 2. Taxonomic overview of the study design for the rise of SLMs in healthcare.
1.4. Contributions
The rise of healthcare SLM is evident from Fig. 1(b). Existing reviews and surveys cover general studies on LLM (Wu et al., 2025c), the application of LLM in healthcare (Guo et al., 2024; Wang and Zhang, 2024), and broad analyses of SLM (Lu et al., 2024a; Van Nguyen et al., 2024; Wang et al., 2024d). Despite the abundance of existing surveys and reviews, as shown in Table 1, there remains a pressing need for a comprehensive examination of the rise of small language models (SLMs) for generative AI applications in healthcare tasks. To address this gap, our survey provides a comprehensive analysis of SLMs, highlighting their optimization strategies, clinical applicability, and transformative potential in resource-limited and safety-critical healthcare settings. To the best of our knowledge, this is the first comprehensive survey on SLMs in healthcare. Our major contributions are as follows:
  1. (1)
    Propose an intuitive taxonomy for tuning and compression strategies in language models, and review prior studies to highlight emerging trends in model development (Fig. 2).
  2. (2)
    Outline the timeline of the SLM development in healthcare to categorize them by parameter range and functionality (Fig. 3).
  3. (3)
    Compile existing open-source resources, including publicly available models (Table 6) and commonly used training datasets (Table 4) in the literature.
  4. (4)
    Examine reductions in carbon footprint and showcase healthcare-specific SLMs to guide researchers in adopting sustainable and efficient practices in the healthcare research domain (Table 5.3).
  5. (5)
    Outline the role of SLMs in healthcare applications (§5.2), examine associated challenges, and highlight key research directions for clinical integration (§6).
We first identified the models through our search strategy, discovered the model configurations, and other architectural foundations. We further investigate its adaptation to clinical precision and compression techniques, followed by outlook as shown in Fig. 2.
Existing studies Year Focus
HS NLP SLM MD D/G
Small language models: survey, measurements, and insights (Lu et al., 2024a) 2024
A survey of small language models (Van Nguyen et al., 2024) 2024
A comprehensive survey of small language models in the era of large… (Wang et al., 2024d) 2024
A survey of personalized large language models… (Liu et al., 2025b) 2025
A survey on LLM-generated text detection (Wu et al., 2025c) 2023
Localizing in-domain adaptation of transformer-based… (Buonocore et al., 2023) 2023
Efficient compressing and tuning methods for large language models… (Kim et al., 2024b) 2023
Large language models in medical and… (Wang and Zhang, 2024) 2024
Large language model for mental health… (Guo et al., 2024) 2024
Our study 2025
Table 1. Focused existing surveys in this field. Here HS: Healthcare-specific, MD: Mental discourse, D/G: Decoder-only or GPT-like.
2. Architectural foundations: Building SLMs from scratch
The lightweight architectures are designed for efficient performance with fewer parameters and reduced computational overhead while maintaining high performance in medical applications. While model efficiency is important, the success of SLMs in healthcare also depends on the quality and accessibility of medical data. This motivates a closer look at the data landscape in clinical settings. In this section, we identify SLMs in healthcare, their timeline, and configurations (§2.1) to focus on architectural foundations in healthcare through healthcare data (§2.2), pre-training components (§2.3), and attention mechanisms (§2.4).
2.1. SLM in healthcare: Timeline and configuration
Figure 3. Timeline of SLM development in healthcare. Transforming NLP-driven clinical decision-making, NLP tasks for medical conditions, and doctor-patient conversation or mental health analysis.
We present the timeline of model development in Fig. 3. SLMs are configured with different aspects to optimize their performance. Notably, few studies provide detailed configurations of SLMs for healthcare applications. Table 2 presents the available configurations we identified, offering an overview of current development efforts. The model size, measured in parameters, typically ranges from 124 million to 7 billion weights. The model depth, defined by the number of transformer layers, typically spans from 22 to 32 layers. The hidden dimension, or the size of the model’s embedding vectors, usually ranges from 1024 to 4096 units. To optimize attention processing, SLMs employ 16 to 64 parallel attention heads, enabling the model to focus on various aspects of the data simultaneously. Additionally, the context window defines the maximum sequence length, allowing the model to process up to 4096 tokens, which is crucial for handling long clinical texts such as patient records and discharge summaries. These design choices ensure that SLMs achieve both high accuracy and efficiency in healthcare applications.
Table 2. Model configurations for existing SLM in healthcare, despite the limited availability in this domain. HS: hidden size, VS: Vocabulary size, MCW: Max context window
Model Name Size Year Attention Layers HS Heads Activation VS MCW
BioGPT 347M, 1.5B 2023 MHA 24 1024 16 GELU 42384 1024
BioMedLM (Bolton et al., 2024) 2.7B 2024 FA 32 2560 20 GELU 28896 1024
BioMistral 7B 2024 FA 32 4096 32 SiLU 32000 4096
mhGPT 1.98B, 2024 - 22 3072 64 - 52000 2048
2.8B 32
2.2. Healthcare data
The healthcare domain presents unique challenges for language models due to the nature of its data. Clinical text data from EHRs, discharge summaries and physician notes holds valuable insights but is underutilized due to unstructured formats and medical jargon (Fu et al., 2023). While models like Med-PaLM (Tu et al., 2024) excel in medical NLP, their scale limits deployment, underscoring the need for efficient and reliable SLM training strategies.
The datasets used for SLM construction are shown in Table 4. Yan et al. (2024) highlights the challenges with data privacy and the lack of the open domain datasets. A potential solution is the synthetic health data generation (SHDG), leveraging learning from statistical distributions of real-world datasets. Recent reviews and surveys on SHDG (Gonzales et al., 2023; Pezoulas et al., 2024) provide a quick reference guide for selecting appropriate synthetic data generation strategies. Recent studies used LLMs for SHDG and validated that LLM-generated synthetic data can improve system performance (Kang et al., 2024). A comprehensive study on metrics for SHDG evaluation (Osorio-Marulanda et al., 2024) noted the absence of a composite metric for quality assessment. SHDG supports data enhancement888augmenting, interpolating, or fusing with synthetic data and de-identification999masking personally identifiable information from patient records in healthcare but also raises concerns about vulnerabilities such as data poisoning attacks101010malicious data is inserted into the training set, intentionally corrupting healthcare models to produce inaccurate or biased outputs that can compromise model integrity. Data poisoning attack (Zhao et al., 2025) occurs when even 0.001% of training tokens are replaced by medical misinformation (Alber et al., 2025). Lobo et al. (2022) emphasizes the need for robust evaluation against data poisoning, which can be mitigated through secure data practices, thorough cleaning, and multistage model training (He et al., 2025).
Model Training Dataset Evaluation Dataset
AlpaCare (Zhang et al., 2023b) MedInstruct-52k MedInstruct-test, iCliniq, MedQA, HeadQA, PubmedQA, MedMCQA, MeQSum, MMLU, BBH, TruthfulQA
Apollo (Wang et al., 2024a) ApolloCorpora (12 multilingual med sources) XMedBench (6 languages)
BioGPT (Luo et al., 2022) PubMed abstracts BC5CDR, KD-DTI, DDI, PubMedQA, BioASQ, HoC
BioMedLM (Bolton et al., 2024) PubMed abstracts and full articles MedMCQA, MedQA, MMLU Medical Genetics, PubMedQA, BioASQ
BioMistral (Labrak et al., 2024) PubMed Central (3B tokens) MMLU-medical, MedQA, MedMCQA, PubMedQA
BioMistral-NLU (Fu et al., 2024) MNLU-Instruct (33 datasets) 15 NLU datasets (BLURB, BLUE)
CancerGPT (Li et al., 2024b) DrugComb (common tissue) DrugComb(Drug pairs in rare tissue)
CancerLLM (Li et al., 2024a) Clinical notes, CancerNER, ICDdiagnosis Phenotype extraction and diagnosis generation; counterfactual and misspelling robustness testbeds
ChatCounsellor (Liu et al., 2023a) Psych8k Synthetic data from interviews
ChatDoctor (Li et al., 2023a) Stanford Alpaca + HealthCareMagic iCliniq, Wikipedia, MedlinePlus
ClinicalMamba (Yang et al., 2024a) MIMIC-III (ICU notes) n2c2 2018, ICD-9 coding
Med-$R^{2}$ (Lu et al., 2025) None (prompt-based only) MedQA, MedMCQA, MMLU-Med
MedAlpaca (Han et al., 2023) Medical Meadow (Anki Flashcards, StackExchange, WikiDoc, CORD-19, PubmedQA, MedQA, OpenAssistant) USMLE
Meditron (Chen et al., 2023) GAP-Replay corpus MedQA, PubMedQA, MMLU-Medical
MentaLlama (Yang et al., 2024b) IMHI dataset (Reddit, Twitter, SMS) Dreaddit, SAD, CAMS, Loneliness, MultiWD, IRF and 4 more
MentalQLM (Shi et al., 2024) DR, MultiWD, Dreddit, Irf, SAD Held-out subsets of DR, MultiWD, Dreddit, Irf, SAD
mhGPT (Kim et al., 2024a) PubMed + Reddit (mental health) IRF, Dreaddit, SAD, MultiWD, PPD-NER
PMC-LLAMA (Wu et al., [n. d.]) MedC-K (papers, textbooks) + MedC-I (instructions) MedQA, MedMCQA, PubMedQA
RadPhi-2 (Ranjit et al., 2024a) Radiopaedia.org, Super-Natural Instructions, Mimic-CXR report data Held-out Radiopaedia QA sets, MIMIC-CXR report data
RadPhi-3 (Ranjit et al., 2024b) Mimic-CXR, MedicalDiff VQA, ChestImagenome, CheXpert Plus reports RaLEs, RadGraph2, MEDNLI, RADNLI, PadChest
Table 3. Comparison of healthcare SLMs: Year, training/evaluation datasets, and data scale.
Task Datasets
Language Inference iCliniq, ICD9 Coding (Mullenbach et al., 2018), i2b2 (Uzuner et al., 2011), MedlinePlus (Li et al., 2023a), MIMIC-CXR (Johnson et al., 2019)
Language Understanding Clinical Knowledge, College Biology, Medical Genetics, Professional Medicine (Hendrycks et al., 2020)
Named Entity Recognition BC5CDR (Li et al., 2016), ICD9 coding (Mullenbach et al., 2018), i2b2 (Uzuner et al., 2011)
Question Answering BioASQ (Tsatsaronis et al., 2015), MedMCQA (Pal et al., 2022), MedQA-USMLE (Kung et al., 2023), PubMedQA (Jin et al., 2019), Radiopaedia QA (Ranjit et al., 2024a)
Relation Extraction BC5CDR (Li et al., 2016), DDI (Herrero-Zazo et al., 2013), KD-DTI (Hou et al., 2022)
Text Classification CAMS (Garg et al., 2022), Dreaddit (Turcan and McKeown, 2019), HoC (Baker et al., 2016), IRF (Garg et al., 2023), Loneliness (Tsakalidis et al., 2019), MultiWD (Sathvik and Garg, 2023), SAD (Mauriello et al., 2021)
Table 4. Open-source datasets used as benchmarks for NLP-driven healthcare tasks in the development of SLMs for clinical applications.
High-quality datasets: The ‘entropy law’ highlights that the compression ratio of the training data is a decisive factor affecting model performance, if the overall quality and consistency of selected samples remain unchanged (Yin et al., 2024). With limited computational resources, the use of high-quality data points is more effective for training than large and noisy datasets, improving efficiency with limited computational resources (Kang et al., 2024). The data cleaning process filters noise, removes duplicates, handles outliers, and balances datasets to ensure better data quality. The discovery of data selection mechanisms ensures the learning of language models from high-quality data, thereby improving efficiency (Zhao et al., 2024). However, poor selection may lead to biased results and reduced model performance (Xie et al., 2023a). A recent survey on data selection mechanism highlights the need to scale down datasets, standardize evaluation metrics and data-centric benchmarking (Albalak et al., 2024). Qin et al. (2024b) discussed data assessment and selection methods to identify the most beneficial data points that are different from standard data quality assessment measures (Lee et al., 2018).
Data cleaning mechanisms are de-identification, de-sensitization, and de-duplication, that enhance model performance by reducing redundancy and increasing diversity. De-duplication improves data quality and prevents data leakage between training and test sets, as demonstrated by models like MediTron (Sallinen et al., 2025) and Apollo (Wang et al., 2024a). Importance resampling efficiently selects relevant data in a reduced feature space, outperforming existing methods and improving downstream model accuracy (Xie et al., 2023a). Notably, Shi et al. (2024) employ perplexity and k-center greedy methods to increase data quality and data diversity, respectively, while developing SLM. After surveying data diversity metrics, Yang et al. (2025) proposed an effective data selection strategy; however, its application in the healthcare domain remains unexplored, underscoring the need to benchmark selection methods for medical datasets to enhance diversity, minimize redundancy, and prevent overfitting. We enlist the training and evaluation datasets used in the existing SLMs for healthcare in Table 3 and mention open source datasets for each NLP task in Table 4.
2.3. Pretraining components
To effectively leverage SLMs, it is essential to understand the underlying components for pre-training the model. Pretraining on diverse, unlabeled corpora enables language models to learn general linguistic patterns, while task-specific fine-tuning allows them to adapt effectively to downstream NLP tasks in healthcare. Following data quality assessment and data selection, SLMs internalize linguistic structures, semantic relationships, and general world knowledge, establishing a robust foundation of language understanding.
Tokenization: Healthcare elements - ICD codes and drug names are challenging due to their ambiguity and domain-specific nature. Word-level tokenization (e.g., Word2Vec) splits text into whole words, leading to a large vocabulary and poor handling of rare words. Character-level tokenization uses individual characters, which handles rare or misspelled words well but results in longer sequences and less linguistic structure. In contrast to both, subword tokenization (e.g., BPE, WordPiece, and the Unigram Language Model) offers a middle ground by breaking words into frequent and meaningful subunits - prefixes or suffixes, mostly used by modern LLMs to efficiently handle both common and rare words (Bradshaw et al., 2025). Tokenization of non-language text (like medical codes) prevents LLMs from recognizing the inherent structure of coding systems, making them less effective at accurate classification (Soroush et al., 2024). Recently, tokenization in clinical texts has improved by using domain-specific tools like abbreviation disambiguation (Hosseini et al., 2024) and medical knowledge bases (e.g., Unified Medical language System (UMLS) (Bodenreider, 2004)) to ensure consistency. Language models used tools like TokenMonster [^11] https://github.com/alasdairforsythe/tokenmonster, a subword tokenizer and vocabulary generator, and BatchBPE (Morgan, 2024), an efficient Python-based Byte-Pair Encoding (BPE) method that reduces memory use by merging token pairs in batches. In general, the choice of tokenizer significantly affects the embedding space of the model and the training costs (Ali et al., 2024). Inspired by efficiency of domain-specific vocabulary and tokenization (Gu et al., 2021), BioMedLM (Bolton et al., 2024) uses a custom BPE tokenizer to prevent the tokenization of healthcare-specific terms (Bolton et al., 2024). Kim et al. (2024a) employed both, pretrained tokenizer and custom tokenizer to develop SLM model for mental health analysis. All evaluated LLMs exhibited poor performance on medical code querying, frequently producing imprecise or fabricated codes, underscoring the need for further research before clinical deployment (Soroush et al., 2024). There is pressing need for specialized tokenization strategies for ICD-codes query.
Vocabulary Size: The vocabulary size is the total number of unique tokens (words, subwords, or characters) that the model can recognize and generate. The larger vocabulary enhance both training efficiency and performance of the LLMs (Takase et al., 2024). In SLMs, a dynamic vocabulary curriculum learning improves pretraining efficiency by adjusting token granularity—using longer tokens for predictable content and shorter ones for complex text—leading to better scaling to vocabulary size (Yu, 2025). Tao et al. (2024) introduced three methods to predict compute-optimal vocabulary sizes for SLMs. However, input and output vocabularies may exhibit distinct scaling behaviors in smaller models, underscoring the need for independent scaling strategies to optimize model design (Huang et al., 2025).
2.4. Attention mechanisms
Attention mechanisms play a central role in enabling language models to selectively focus on relevant information within input sequences. In this section, we describe core attention architectures and efficient attention mechanisms along with their time complexities and discuss recent efficiency-driven adaptations designed to support SLM development.
Core attention architectures: Core attention mechanisms in language models can be optimized for healthcare by investigating how clinical concepts and contextual information interact during processing. The attention mechanism has evolved from self-attention to cross-attention, and further to multi-head, entity-aware, and knowledge-guided variants — each contributing to advancements in healthcare applications. For example, Self-attention is useful in attending certain medical aspect— symptoms, diagnoses, or medications, capturing dependencies like the relationship between “chest pain" and “ECG findings" (Vaswani et al., 2017). Cross-attention enables one source of information (e.g., lab results) to attend to another (e.g., physician notes), supporting rich understanding by integrating information from both structured and unstructured information (Lai et al., 2024). Global attention highlights key clinical anchors (e.g., “principal diagnosis” or “chief complaint”) that influence interpretation of the entire record (Liu, 2025). Multi-head attention (MHA) models multiple types of clinical relationships in parallel-temporal patterns, medication-disease interactions, or symptom clusters (Meng et al., 2025). The knowledge-guided attention integrates biomedical ontologies or knowledge graphs (e.g., UMLS) to bias attention toward medically validated relationships, enhancing clinical accuracy and explainability (Li et al., 2023b). However, a key limitation of core attention mechanisms—particularly self-attention—is their memory inefficiency, as the computational and memory cost scales quadratically with the input length, making it challenging to process long clinical documents or patient histories in EHR.
Efficient attention: To address computational bottlenecks, SLM in healthcare leverage memory-efficient attention adaptations. Sparse attention restricts these interactions to the most relevant information— focusing only on recent or abnormal findings—making it ideal for processing lengthy medical notes efficiently (Beltagy et al., 2020; Child et al., 2019). Multi-query attention (MQA) allows multiple attention heads to share the same keys and values (KV) pairs (Xu et al., 2023). For example, in a clinical note, different attention heads can be used to extract aspects like demographic information, medication history, and clinical conditions, while sharing the same KV pairs. Grouped-query attention (GQA) groups query heads into clusters, with each cluster sharing a KV pair (Ainslie et al., 2023). For example, when extracting SDoH, one cluster of heads might focus on housing, another on employment, and another on education. Multi-Head Latent Attention uses low-rank compression of KV pairs to reduce the need for a KV cache (Lu et al., 2024a). For example, a model extracting QoL measures from patient feedback can use this compression to handle large datasets without storing full-dimensional attention matrices for each concept. QCQA introduces a quality- and capacity-aware grouped query attention mechanism that optimally balances key-value (KV)-cache size and text generation accuracy, outperforming GQA by up to 20% in accuracy and reducing memory requirements by 40% without fine-tuning (Joshi et al., 2024). Flash Attention (FA) divides the attention matrix into smaller blocks and recomputes them during backpropagation, saving up to 4 times more memory without losing precision (Dao et al., 2022). For example, in a clinical question-answering system, instead of storing the entire attention matrix for a patient’s full medical report, FA stores only key blocks and recomputes others as needed, reducing memory usage and allowing the system to handle larger datasets like multi-visit histories.
Low-Rank Factorization approximates the full attention matrix by decomposing it into smaller, lower-dimensional matrices, retaining only the essential relationships. In a clinical concept extraction model, this approach increases efficiency by replacing large KV matrices with more compact representations. A series of advancements in low-rank adaptation techniques have emerged to optimize LLM fine-tuning. Memory-Efficient Attention optimizes processing by calculating attention scores without storing the entire matrix. In a real-time clinical decision support system, this approach allows for dynamic attention score calculation using medical records, enabling efficient analysis of patient symptoms and medical history without storing full context. Sliding Window Attention captures local dependencies by focusing on a fixed-sized window of neighboring tokens. The clinical note SUM allows the model to focus on specific sections such as the history of the present illness. Random Projection Attention uses dimensionality reduction to approximate softmax attention, improving scalability. Building on these foundations, recent advancements have given rise to more sophisticated architectures. Notably, Mamba (Gu and Dao, 2023) introduces a selective state space model characterized by input-dependent transitions. In predictive models for patient outcomes, it reduces the complexity of processing high-dimensional data, such as medical tests and demographics, making the model more efficient and computationally cost-effective for large datasets and real-time predictions. We analyze the time complexity of each attention mechanism to shed some light on how healthcare research can integrate attention-driven component based on available resources in Table 5.
Table 5. Time complexities of various attention mechanisms in clinical NLP, from least to most efficient. Here, $N$ is the sequence length, $d$ is the embedding dimension, $M$ is the cross-attended length, $w$ is a local window size, $g$ is the number of global tokens, and $r$ is a low-rank dimension. “KG overhead” and $\epsilon$ indicate extra constant factors due to knowledge bases or entity-related modules. These values are based on official reports or estimated according to functionality.
Attention type Time complexity Use case in healthcare
Self-Attention $O(N^{2}\cdot d)$ Intra-note relationships
Cross-Attention $O(N\cdot M\cdot d)$ Structured + unstructured fusion
Global (Sparse) Attention $O(N\cdot w+N\cdot g)$ Global tokens for long notes
Multi-Head Attention $O(N^{2}\cdot d)$ Parallel clinical patterns
Knowledge-Guided Attention $O(N^{2})+\text{KG overhead}$ Ontology-driven alignment
Sparse Attention $\sim O(N\cdot w),\,O(N\log N),\dots$ Restrict to local windows / relevant tokens
Multi-Query Attention $O(N^{2}\cdot d)$ [reduced memory] Heads share the same K/V pairs
Grouped-Query Attention (GQA) $O(N^{2}\cdot d)$ [fewer K/V sets] Group queries into clusters
Flash Attention $O(N^{2}\cdot d)$ [blockwise] Near full precision, reduced memory
Multi-Head Latent Attention $O(N^{2}\cdot r),\;(r\ll d)$ Low-rank latent representations
Low rank factorization $O(N^{2}\cdot r),\;(r\ll d)$ Dimensionality reduction for large clinical corpora
Memory efficient attention $O(N^{2}\cdot d)$ Handle lengthy EHR data on limited GPU/TPU resources
Single window attention $O(N\cdot w)$ Focus on a single local segment (useful for chunk-based analysis)
Random Projection Attention $O(N)$ Linear approximation of full attention
3. Prompting, structured fine-tuning and reasoning: Adapting SLMs to clinical precision
After pretraining and the integration of attention mechanisms, prompt engineering, reasoning and instruction tuning becomes the most essential components to adapt the general instruct-language model to specific healthcare tasks. Fine-tuning enables the model to align with domain-specific objectives and improve performance in clinical settings. In this section, we discuss in-context learning (§3.1) and fine-tuning (§3.2).
3.1. In-context learning
In-context learning (ICL) in healthcare is the ability of language models to learn and adapt to new healthcare NLP tasks and provide meaningful outputs based on the context in clinical notes, without requiring explicit retraining or fine-tuning. While Dong et al. (2022) provides a broad overview of ICL, a comprehensive survey dedicated to its applications in cNLP remains absent. In-context learning enables models to adapt to tasks without gradient updates, while preference alignment further refines their behavior to align with human values and domain-specific expectations, ensuring safer and more relevant responses in healthcare.
Alignment: The most renowned method for preference alignment is reinforcement learning from human feedback (RLHF), which relies on reward models and policy updates that often introduce instability due to probability ratio-based optimization (Bai et al., 2022). DeepSeek’s Group Relative Policy Optimization (GRPO) made RLHF more efficient (Shao et al., 2024), especially for complex reasoning tasks, while Stanford’s direct preference optimization (DPO) refines clinician response selection by dynamically adjusting the log probabilities of preferred versus non-preferred responses. These models uses human feedback to train language models and generate more preferred responses without requiring a separate reward model. Past studies reported that supervised fine-tuning (SFT) is sufficient for healthcare CLS tasks, whereas direct performance optimization yields superior results for clinical reasoning, and clinical triage (Savage et al., 2024). Other than DPO (Ahn et al., 2024), the research community has explored multi-agent approach of preferential data annotation through multiple rounds of dialogue (Zhang et al., 2025a; Gururajan et al., 2024) and reinforcement learning with environmental feedback (Liu et al., 2024b) in healthcare applications. Thus, preferential alignment methods have the potential to alleviate the labor-intensive workload of clinicians.
Prompt engineering: Prompt Engineering is the process of designing and optimizing inputs (prompts) given to the SLMs to generate meaningful, relevant, and actionable outputs in the healthcare domain (Liu et al., 2023c). Recently, the research community reported comprehensive reviews and surveys on the latest advances in prompt engineering for healthcare (Wang et al., 2025b). As the choice of prompt can notably influence the accuracy of LLM responses to medical queries (Wang et al., 2024b), Prompt engineering has emerged as a critical skill for healthcare professionals, enabling more effective interaction with language models (Meskó, 2023). Building on this, frameworks like MED-Prompt have been developed to construct clinically relevant prompts that guide models in interpreting and extracting information from medical texts (Ahmed et al., 2024). More recently, the introduction of active prompting—an adaptive technique where the model iteratively refines its responses based on prior feedback—has shown promise in enhancing diagnostic reasoning and minimizing erroneous inferences (Diao et al., 2023). A recent survey by Wang et al. (2023a) reported a range of methods for designing prompts in healthcare applications.
Reasoning: Reasoning further increases the capabilities of model to infer relationships between concepts, drawing conclusions, and applying domain knowledge to produce outputs that are logically consistent with medical standards. Huang and Chang (2022) discuss different mechanisms for improving and eliciting reasoning in LLMs and benchmarks methods for evaluating reasoning abilities. To encourage reasoning in LLMs, chain-of-thought (CoT) prompting (Wei et al., 2022) and “let us think step-by-step" prompting (Kojima et al., 2022) guides them to generate explicit reasoning steps before answering. CoT allows structured inference by breaking down complex medical queries into intermediate reasoning steps in healthcare NLP tasks (Wang et al., 2024e; Jiang et al., 2025). Wu et al. (2025a) critiques CoT prompting for its limitations across subspecialties and k-shot approaches for introducing noise, proposing AutoMedPrompt, leveraging textual gradients for optimizing system prompts to enhance medical reasoning in foundation LLMs.
The integration of external knowledge has advanced healthcare informatics by enhancing factual accuracy and reducing hallucinations (Li et al., 2025; Wang et al., 2024c). Retrieval-augmented generation (RAG) further strengthens healthcare language models by incorporating clinical knowledge into text generation to deliver contextually relevant responses (Lahiri and Hu, 2024). The integration of RAG and CoT demonstrates improved performance on high-quality clinical notes, where early retrieval effectively anchors subsequent reasoning steps in domain-specific evidence while CoT-driven RAG offers advantages in handling lengthy and noisy notes by guiding retrieval through structured reasoning (Wu et al., 2025b; Li et al., 2024e). In MetaICL (2023), meta-learns task-agnostic prompts for generalizing across specialties (e.g., cardiology to nephrology), “learning" patterns dynamically from the context (Min et al., 2021). The integration of structured knowledge from a knowledge graph (KG) into the model’s training or inference process enhance the model’s ability to understand context and reason over relationships. KG has the potential to map patient data to structured ontologies (e.g., SNOMED-CT, UMLS) and infer relationships (e.g., "fatigue + jaundice → hepatitis risk"). Recent studies on neuro-symbolic AI suggest the need of further investigation of metacognition121212model’s capacity to monitor, assess, and adapt its own reasoning or output generation processes, often by estimating confidence, recognizing uncertainty, or reflecting on errors. for advancing the field towards more intelligent, reliable, and context-aware AI systems (Colelough and Regli, 2025; Didolkar et al., 2024). Metacognition improves decision-making in healthcare practitioners through cognitive bias identification, facilitation of systematic approaches, holistic approaches, and creative thinking (Church and Carroll, 2023). MD-PIE emulates cognitive and reasoning abilities in medical reasoning and decision-making (Esteitieh et al., 2025). However, more robust evaluation frameworks that incorporate metacognitive abilities are required as they are essential for developing reliable language model enhanced clinical decision support systems (Griot et al., 2025).
3.2. Supervised fine-tuning
Fine-tuning on smaller, task-specific datasets enables language models to effectively leverage pre-trained knowledge for specialized domains or tasks (Fu et al., 2024; Tran et al., 2024). These techniques are address challenges like limited computing resources, data quality, availability, and robustness, ensuring efficient adaptation to new tasks without extensive retraining. Instruction tuning is sometimes used interchangeably with instruction fine-tuning; however, the latter often refers to training with chat-specific templates. While the distinction is debated in the literature, we do not focus on this differentiation in this work.
Instruction tuning: Instructing tuning leverages instructions for flexibility. Instruction datasets are created by integrating annotated data into instruction-output pairs using templates. Alternatively, LLMs can generate outputs from manually or seed-generated instructions. In the human-generated questions with LLM annotations approach, humans create diverse, relevant questions, while LLMs provide initial answers. These answers are then reviewed and refined by humans, resulting in high-quality datasets for instruction tuning. The LLM-generated SDHG for instruction tuning increases training example diversity and reduces manual effort. Nazar et al. (2025) provides A detailed guidance on designing and evaluating instruction datasets for healthcare is available appropriate data construction methods and strategies ensuring data quality (Zhang et al., 2023a; Nazar et al., 2025). BioInstruct (Tran et al., 2024) and MIMIC-Instr (Wu et al., 2024) are the high-quality healthcare instruction datasets designed to enhance the training of language models for clinical NLP tasks.
Instruction fine-tuning Instruction fine-tuning leverages focused instructions for specific tasks. A recent study suggests that while full parameter instruction fine-tuning can enhance task-specific performance, it may also lead to knowledge degradation and an increased risk of hallucination, potentially due to the overwriting of pre-trained knowledge during the fine-tuning process (Ghosh et al., 2024). AlpaCare is fine-tuned on MedInstruct-52k, a diverse machine-generated medical instruction-finetuning dataset, that outperforms previous medical LLMs (Zhang et al., 2023b). LlamaCare enhances LLMs’ ability to predict outcomes by fine-tuning existing LLM on various clinical tasks (Li et al., 2024c). Instruction fine-tuning is employed to adapt the SLM for biomedical tasks (Luo et al., 2024), medical entity recognition (Liu et al., 2024c), radiology tasks (Ranjit et al., 2024a), ICU readmissions (Namvar et al., 2025), medical documentation (Leong et al., 2024) and billing and coding (Rollman et al., 2025). However, the high-quality, diverse, and relevant instructions are key for SLM fine-tuning.
Parameter efficient fine tuning (PEFT) adapts models to specific tasks by training only a small (upto <5%) subset of parameters making them deployable on edge devices such as tablets in rural clinics. This rapid customization fine-tunes the model more efficiently while preserving the advantages of the pre-trained model. Starting with LoRA, subsequent methods like LoRA+, LoRA-FA, and AdaLoRA improve efficiency, memory usage, and adaptability, while approaches such as LoRA-drop, and DoRA introduce novel strategies like random matrix adaptation, output-based pruning, and weight decomposition to further enhance parameter-efficient tuning (Mao et al., 2025). LoRA injects low-rank matrices into existing weights, allowing models to adapt to specialized tasks without full retraining, preserving foundational medical knowledge with minimal overhead (Shi et al., 2024). Recent fine-tuning of Phi3.5-mini [^13] https://huggingface.co/microsoft/Phi-3-mini-4k-instruct through quantized low-rank adapters (QLoRA) underscores its potential for providing basic medical recommendations in telemedicine and chatbot systems. Adapter modules, inserted between transformer layers, facilitate multi-task learning on these small layers instead of modifying the base model as mixture of experts (Cai et al., 2025). Domain-specific PEFT approach (Gajulamandyam et al., 2025) is employed for hospital discharge paper SUM (Goswami et al., 2024), clinical RE (Zhou et al., 2023a), and biomedical NER (Abishek et al., 2024). Prompt tuning optimizes continuous “soft prompts" to guide models toward structured outputs (Wang et al., 2023c; He et al., 2024). In prefix tuning a small set of trainable parameters (prefixes) is prepended to the input representation in a frozen language model, allowing adaptation to domain-specific tasks without modifying the core model weights.
4. Compression techniques: Ensuring accessibility and sustainability
The need for compressed models—via quantization, pruning, and knowledge distillation—arises from the demand to deploy LLMs efficiently on resource-constrained devices, reduce inference latency and memory usage, cut energy costs, and enable real-time applications while maintaining performance close to their larger counterparts. In healthcare, compressed models enable real-time clinical decision support, preserving patient privacy while operating on limited hardware for NLP-driven clinical decision-making.
4.1. Knowledge distillation
Knowledge distillation (KD) is a model compression technique that trains a smaller student model to replicate the behavior of a larger teacher model (Hinton et al., 2015). KD has proven effective for creating high-performance SLMs from LLMs while significantly reducing model size (Wang et al., 2024d). KD is comprised of two components: knowledge and distillation. The knowledge covers a variety of teacher-student frameworks: labeling (teacher-generated outputs), expansion (in-context sample generation), data curation (meta-driven synthesis), feature extraction (teacher-derived internal representations), feedback (evaluative guidance on student outputs), and self-knowledge (student-led output generation and quality filtering) (Xu et al., 2024). Distillation covers supervised fine-tuning, divergence and similarity, Reinforcement learning, rank optimization. In supervised fine-tuning, also known as sequence-level KD, the student model is trained to maximize the likelihood of teacher-generated sequences, aligning its predictions with the teacher’s outputs through divergence or similarity-based loss functions. Similarity-based methods aligns internal representations of teacher and student models while divergence-based methods minimize distributional differences. Reinforcement learning-based distillation leverages teacher-generated feedback to train reward models and optimize student policies, aligning student outputs with teacher preferences through reward maximization and divergence minimization. Ranking optimization offers a stable, sample-efficient alternative to reinforcement learning by directly fine-tuning language models to favor preferred outputs over less favored ones using fixed teacher-generated preference datasets. Furthermore, the KD methods are categorized into white-box vs. black-box based on the information accessible from the teacher (Yang et al., 2024c).
In white-box KD, a student not only observes the teacher’s outputs but also has access to the teacher model’s internal parameters or representations (Agarwal et al., 2024). This additional insight into the teacher’s “internal structure" often enables the student to perform better since it can directly leverage the teacher’s learned features and hidden knowledge, addressing the challenges of LLM scalability and portability for health event prediction (Ding et al., 2024). A feature-level knowledge distillation technique was introduced for medication recommendation, thereby training the SLM using hidden states of LLM (Liu et al., 2024d). In black-box KD, by contrast, only the teacher’s outputs (e.g. predicted logits or generated responses) are available to the student; the “internal structure" of the teacher is treated as a black box (Peng et al., 2023; Wang et al., 2023b). This scenario commonly arises when using proprietary LLM APIs as teacher models, where only input–output pairs are accessible, but the underlying model weights remain unavailable. Despite the limited information, black-box KD has shown promising results by fine-tuning students on the input–output behaviors of large models. Complex reasoning and problem-solving skills can be transferred by distilling the teacher’s CoT and prompt-based capabilities.
Discussion on KD: Both black box and white box KD mechanisms are widely used in healthcare for developing SLM. Some notable studies are employing white-box KD for medication recommendation (Liu et al., 2024d), health event prediction, lung cancer detection (Pavel et al., 2024), emotion recognition in daily healthcare (Zhang et al., 2024a), clinical IE (Vedula et al., 2024) and medical data SUM (Liu et al., 2024a). A sleep health management SLM leverages few-shot CoT (black-box) distillation to transfer problem-solving strategies, long-tail expert knowledge, and personalized recommendation capabilities from LLMs into the SLM  (Zheng et al., 2024). Key problems identified while deploying knowledge distillation in the medical domain are insufficient training data, lack of data sharing, long-tail data distribution, label imbalance, and potential for further performance improvement (Meng et al., 2021). In Adapt-and-Distill, both the teacher (LLM) and the student (SLM) are first adapted or fine-tuned to the target domain or task distribution, followed by knowledge distillation (Yao et al., 2021). A task-agnostic, domain-adaptive distillation approach creates small, fast, and effective models by expanding domain-specific vocabulary and compressing large pre-trained models.
4.2. Pruning
Pruning is a long-established technique for model compression that reduces the size of a model by removing unnecessary or less important parameters (LeCun et al., 1989; Han et al., 2015). In language models, many weights can be zeroed out or eliminated with little to no effect on overall performance (Bergsma et al., 2025). By pruning these redundant parameters, we can significantly shrink model size and accelerate inference, with only a minor drop in accuracy if done carefully. Pruning thus makes models more storage-friendly, memory-efficient, and computation-efficient. Based on the granularity of removal, pruning is categorized as unstructured pruning and structured pruning.
Structured pruning streamlines LLM by eliminating entire architectural components—such as neurons, filter channels, attention heads, or even layers—instead of removing individual parameters (Ma et al., 2023; Xia et al., 2023). By selectively pruning structurally coherent groups of parameters, this approach yields a smaller yet regularly structured model—maintaining dense connectivity while reducing inference costs, for instance, LLM-Pruner (Ma et al., 2023). Unstructured pruning removes individual weights from LLM without considering any specific structure within the model (Frantar and Alistarh, 2023). Unstructured pruning removes less important weights based on a threshold, resulting in a sparse model where the architecture remains unchanged, as long as enough weights stay to maintain performance. Unstructured pruning can deliver higher accuracy for a given compression ratio by allowing more targeted weight removal. However, its irregular sparsity pattern leads to inefficient memory access and limited speed-ups on standard hardware. To fully leverage the compression benefits, specialized libraries or hardware are needed, and even then, acceleration gains are modest unless the model is highly sparse. Post unstructured pruning, significant retraining or fine-tuning is often required to recover lost accuracy that can be resource-intensive.
Discussion on pruning: Contextual pruning is a promising method for building domain-specific language models (Valicenti et al., 2023). Recent research has focused on one-shot pruning methods and improved criteria to minimize the need for costly retraining (Zhang et al., 2024b), validating its performance for healthcare tasks. ATP (All-in-One Tuning and Structural Pruning) is a unified approach that dynamically identifies optimal substructures during fine-tuning, integrating structural pruning with LoRA-aware regularization, achieving superior performance recovery compared to two-stage pruning methods in domain-specific applications (Lu et al., 2024b). The research community has witnessed a range of pruning mechanisms for generic language models (Kim et al., 2024b), however, its investigation for healthcare domain remains unexplored.
4.3. Quantization
Quantization is the process of reducing precision of the model’s weights and activations from floating-point values (e.g., 32-bit) to lower-bit representations (e.g., 8-bit), thereby drastically reducing the memory footprint and computational cost of the model (Gray and Neuhoff, 1998). Quantization reduces model size by up to 4x (for 8-bit) and accelerates inference by using efficient low-bit operations. While it may cause slight precision loss due to mapping floating-point values to discrete levels, careful techniques can minimize accuracy degradation. Three components for quantization are weight, activation, and fixed-point (Bai et al., 2024). Quantization is categorized based on the timing of quantization in the model’s lifecycle: Post-Training Quantization (PTQ) and Quantization-Aware Training (QAT).
Post-training Quantization (PTQ) compresses a model after it has been fully trained by quantizing its weights (and optionally activations) to a lower precision, without requiring any additional retraining (Dettmers et al., 2022). Post-Training Quantization (PTQ) is a simple and efficient method where an already trained LLM is converted to lower precision (e.g., 8-bit) with minimal data, without requiring full fine-tuning. It is faster and more cost-effective than retraining but may compromise on accuracy due to quantization errors. In practice, PTQ methods use an engineering approach to reduce accuracy loss. Quantization-Aware Training (QAT) simulates low-bit precision during training or fine-tuning process, allowing the model to adapt to learn and adjust for quantization errors (Liu et al., 2023b). This typically leads to better accuracy than PTQ, especially at very low bit widths. Mixed precision training in SLMs use both 32-bit (FP32) and 16-bit (FP16 or BFloat16) floating-point formats to optimize performance. Mixed precision allows researchers to train and fine-tune models more efficiently on limited hardware, making it a vital technique for resource-constrained environments and small-scale deployments.
Discussion on quantization: DFQ-SAM is a data-free PTQ approach that calibrates quantization parameters without using original data, helping protect healthcare data privacy during model compression (Li et al., 2024d). Quantized LLM were noticed for virtual mental health assistant (Kumar et al., 2024), and medical documentation (Goyal et al., 2024). Bolton et al. (2024) computes precision with low precision but storage and gradient communication with higher precision. A recent study examines two key methods of quantization: Activation-aware Weight Quantization (AWQ), which selectively preserves important weights to maintain performance, and BitsandBytes (BnB), which applies uniform fixed-precision quantization for constructing BioMistral (Labrak et al., 2024). QLoRA-based quantization enables low-precision LLMs to maintain strong reasoning and classification performance, enabling efficient deployment on edge devices for mental health applications (Shi et al., 2024).
5. SLM in healthcare: Outlook
Healthcare professionals expect accurate answers and proactive guidance in articulating their concerns, correct misunderstandings, and ask clarifying questions—achieved through multi-round conversation training to enhance response credibility. Multi-round training is feasible for SLMs than LLMs due to low computational demands, supporting an efficient implementation of on-device reasoning frameworks (e.g., RAG), enabling real-time personalized medical responses. A unified evaluation platform for healthcare SLMs should incorporate diverse benchmark NLP tasks, a well-defined care continuum (diagnosis, treatment), and comprehensive metrics (confidence, safety and factuality) to enable fair comparisons and advancements in the field. Healthcare-specific SLMs are commonly developed using one of three model sources: (i) generic LLMs, (ii) healthcare-specific LLMs, and (iii) generic SLMs. The generic LLMs are pre-trained generic data, serving as a robust backbone for downstream healthcare applications due to their strong linguistic, commonsense and reasoning capabilities. These models are often optimized and fine-tuned on medical corpora for cNLP tasks. Second, healthcare-specific LLMs are pre-trained on medical texts, explicitly designed to understand clinical jargon, biomedical terminology, and the patient-physician communication. The knowledge is further compressed into smaller architectures to form task-optimized SLMs. Third, generic SLMs are inherently lightweight and serve as a compact and computationally efficient foundation. Their adaption to healthcare domain through tuning, aids hospital systems with strict latency and privacy requirements. This section discusses the current state-of-the-art (§5.1), the role of SLM in healthcare (§5.2), and assessment models (§5.3).
5.1. Current state-of-the-art
We assemble the evaluation of existing healthcare SLMs on various NLP tasks, comparing their performance in medical QA through accuracy in Fig. 4(a), NLU through F-score in Fig. 4(b), and mental health analysis through weighted F-measure in Fig. 4(c). We do not claim any computational novelty, but carry over the results from existing studies to report the current state-of-the-art.
(a) For medical question answering.
(b) Medical language understanding
(c) Mental health analysis
Figure 4. Compiled performance of healthcare SLMs for the most commonly leveraged datasets in healthcare-specific NLP tasks.
Medical question answering: Among all others, the results of six SLMs/versions of healthcare-related models developed for medical tasks (BioMedLM (Bolton et al., 2024), BioMistral-SLERP (Labrak et al., 2024), BioMistral-DARE (Labrak et al., 2024), MediTron (Chen et al., 2023), Med-$R^{2}$ (Lu et al., 2025), PMC-LLAMA (Wu et al., [n. d.])) were comparable on three most commonly used medical QA datasets: PubMedQA (Jin et al., 2019), MedMCQA (Pal et al., 2022), and MedQA-USMLE (Kung et al., 2023). In addition to the medical tasks, we collected results from conversational models in healthcare (ChatDoctor (Li et al., 2023a), MedAlpaca (Han et al., 2023), and AlpaCare (Zhang et al., 2023b)) Fig. 4(a) shows no single model dominates across all medical QA datasets. PMC-LLAMA achieves the highest score on PubMedQA (77.09%), demonstrating strong factual recall in biomedical literature. On MedQA-USMLE, which emphasizes clinical reasoning, Med-$R^{2}$ significantly outperforms all others with a score of 81.06%, indicating its robustness in handling complex medical exam-style questions. BioMedLM shows the most consistent performance across all datasets, ranking among the top models for MedMCQA (57.3%), a multiple-choice QA dataset focused on general medical knowledge. These results underscore the importance of task-specific fine-tuning and diverse pretraining corpora in building high-performing medical language models. The effectiveness of healthcare SLMs varies across datasets, reflecting differences in reasoning complexity and domain specificity. The lack of a universally dominant model highlights the need for adaptive SLM architectures for clinical settings. Incorporating modular training strategies or task-specific adapters could help bridge the performance gap across diverse datasets.
Natural language understanding (NLU): In Fig. 4(b), we collect results for four model (BioMedLM (Bolton et al., 2024), BioMistral-SLERP (Labrak et al., 2024), BioMistral-DARE (Labrak et al., 2024), MediTron (Chen et al., 2023),) evaluated on the four common domains of NLU score: clinical knowledge, professional medicine, college biology, and medical genetics (Hendrycks et al., 2020). BioMistral-DARE consistently achieves the highest scores across all domains, including Clinical Knowledge (63.1%), Professional Medicine (57.4%), College Biology (63.4%), and Medical Genetics (63.3%), highlighting its effectiveness in both foundational and applied medical contexts. BioMistral-SLERP closely follows with its strongest result in Medical Genetics (69%), suggesting robust performance in specialized biological domains. In contrast, MediTron lags across all domains, indicating limited domain adaptation or insufficient training exposure to complex medical content. Fig. 4(b) highlights the varying effectiveness of models, suggesting the need for SLMs with improved efficiency for hospital utilization.
Mental health analysis: We gather the results from three models: MentaLLAMA (Yang et al., 2024b), mhGPT (Kim et al., 2024a), and MentalQLM (Shi et al., 2024); among which MentaLLAMA have two variants: base model and chat model. The more widely adapted datasets are interpersonal risk factors (Garg et al., 2023), Dreddit (Turcan and McKeown, 2019), SAD (Mauriello et al., 2021), MultiWD (Garg et al., 2024), and DR (Pirina and Çöltekin, 2018). MentalQLM achieves the highest scores on IRF (76.81%), Dreddit (82.85%), SAD (67.06%), and MultiWD (77.88%), reflecting its robustness in identifying psychological distress, sentiment, and well-being markers as shown in Fig. 4(c). For the DR dataset, which lacks a score for mhGPT, MentalQLM narrowly outperforms other models with a top score of 84.20%. MentaLLAMA-chat-7B also shows competitive performance, particularly in DR (83.95%) and MultiWD (75.79%), indicating its strength in conversational contexts. The superior performance by existing SLMs underscores the importance of contextual and sensitive language understanding in mental health applications. ChatCounselor is benchmarked against existing LLMs using a robust GPT-4–based evaluation protocol, which emphasizes natural, human-like dialogue over mechanistic response listing (Liu et al., 2023a). ChatCounselor outperforms most open-source models and approaches ChatGPT in quality, underscoring the impact of domain-specific fine-tuning on enhancing personalized mental health support. Future work should explore benchmarking clinical mental health SLMs against existing mental health–focused or general-purpose LLMs to strengthen the confidence of key stakeholders- healthcare professionals. Moreover, as the training datasets are primarily derived from social media, the development of a well-established model is essential for effective deployment in clinical or hospital settings.
Miscellaneous: We further observed three biomedical language models—BioMistral, BioMistral-NLU, and BioGPT—on the RE task using the F1 score as the evaluation metric. BioGPT outperforms the other healthcare SLMs. The ClinicalMamba 2.8B achieves an F1 score of 56.51 and 74.34 for IE task task of rare and common ICD coding, respectively. The performance of various language models on NER tasks across six biomedical datasets demonstrates superior performance by BioMistral-NLU. However, the F1 score varies from 55% to 90% for different NER dataset, highlight the need of specialized medical NER models. Clinical conditions (e.g., cancer) and medical expertise (e.g., radiology) consider the construction of specialized models using diagnosis generation tasks/RE, and radiology NLI/IE, respectively. The current state of the art highlights the need for language models that achieve an optimal balance between computational efficiency and high predictive accuracy, while ensuring responsible use of resources.
Paper Year Base Model #Param ICL FT KD PR QT Task A
BioGPT (Luo et al., 2022) 2022 GPT-2 347M, 1.5B RE, QA, CLS
BioMedLM (Bolton et al., 2024) 2024 GPT-2 2.7B QA, IE, NLU
CancerGPT (Li et al., 2024b) 2024 GPT-2 124M Drug synergy prediction
ClinicalMamba (Yang et al., 2024a) 2024 Mamba 130M Cohort selection, ICD coding
MentalQLM (Shi et al., 2024) 2024 Qwen1.5 0.5B Mental health classification
mhGPT (Kim et al., 2024a) 2024 GPT-NeoX 1.98B, 2.8B Mental health classification
RadPhi-2 (Ranjit et al., 2024a) 2024 Phi-2 2.7B Radiology QA, CXR NLP
RadPhi-3 (Ranjit et al., 2024b) 2024 Phi-3 3.8B summarization, labeling
AlpaCare (Zhang et al., 2023b) 2023 LLAMA-2 7B QA, NLU
Apollo (Wang et al., 2024a) 2024 Qwen <=7B Multilingual Med QA
BioMistral (Labrak et al., 2024) 2024 Mistral 7B QA, multilingual med data
BioMistral-NLU (Fu et al., 2024) 2024 Mistral 7B NER, RE, DC, QA, NLI
CancerLLM (Li et al., 2024a) 2024 Mistral-7B 7B Phenotyping and text gen.
ChatCounsellor (Liu et al., 2023a) 2023 Vicuna-v1.3 7B QA from hospital interviews
ChatDoctor (Li et al., 2023a) 2023 LLaMA-2 7B Dialogue, QA, Retrieval
Med-$R^{2}$ (Lu et al., 2025) 2025 Qwen2.5 7B Retrieval + reasoning pipeline
MedAlpaca (Han et al., 2023) 2023 LLAMA 7B QA and Summarization
MentaLlama (Yang et al., 2024b) 2024 LLaMA-2 7B Mental health generation
Meditron (Chen et al., 2023) 2023 LLaMA-2 7B Med QA, reasoning
PMC-LLAMA (Wu et al., [n. d.]) 2023 LLaMA-2 7B Medical QA, dialogue
Table 6. Comparison of existing SLM in healthcare: Here, ICL: In-context learning, FT: Fine tuning, KD: Knowledge distillation, PR: Pruning, QT: Quantization. The tasks are RE: Relation extraction, CLS: Text Classification, QA: Question Answering, IE: Information extraction. A is availability. indicates that code is not released as on 1st April 2025.
Discussion on comparative analysis of SLMs: SLMs have rapidly evolved for a diverse range of healthcare applications. Table 6 summarizes existing healthcare SLMs, detailing their base architectures, adaptation techniques, and target NLP tasks. Most SLMs are built on lightweight architectures like GPT-2, Phi-2, Mamba, Qwen, and LLaMA-2, often adapted with techniques such as fine-tuning and in-context learning. However, fewer models leverage quantization. Models like BioMistral, BioMistral-NLU, MentalQLM, and PMC-LLAMA exemplify state-of-the-art performance through multi-task adaptability, multilingual benchmarking, and specialized reasoning tasks (e.g., CoT QA). Yet, many efforts are still concentrated on QA and CLS, while areas like clinical SUM, temporal reasoning, and wellness dimension extraction remain underexplored. Additionally, we mapped the availability of SLMs within healthcare domain.
At present, healthcare models employ knowledge-aware methods, including ontology grounding using UMLS or SNOMED CT and RAG to enhance factual correctness. Integrating advanced interpretable reasoning mechanisms, such as metacognition, holds promise for enhancing the trustworthiness and reliability of clinical decision support systems. There is a pressing need to broaden the scope of tasks beyond traditional disease-centric applications like question answering and classification, toward holistic patient care, including the extraction and interpretation of emotional, social, behavioral, and wellness dimensions. Establishing standardized benchmarks for emerging clinical NLP tasks - wellness assessment, mental health triage, and personalized care planning will be vital to evaluating real-world utility and scientific progress. Moreover, Magnini et al. (2025) evaluates the use of locally deployed open-source generic SLMs for medical chatbots, focusing on hypertension self-management to ensure privacy and reliability. It demonstrates that lightweight SLMs can perform competitively in key tasks, offering a viable, privacy-preserving alternative to large proprietary models.
5.2. The role of SLM in healthcare
Beyond the performance of standalone models, we explore the role of SLMs in synergy with LLMs, emphasizing their complementary strengths, capabilities, and applications in healthcare research. The potential of collaborative and competitive roles of SLMs and LLMs is underappreciated in fostering efficient and accessible AI models for clinical decision making. Amid the broad spectrum of SLM applications observed in prior studies on general models (Chen and Varoquaux, 2024), this section delineates their specific relevance to NLP-driven decision-making tasks within the scope of healthcare informatics.
(a) Data curation
(b) On device personalization
(c) Ensembling
(d) Distillation
Figure 5. The role of SLM in healthcare
Data curation: Zhou et al. (2023b) proved that fine-tuning with just 1,000 high-quality instruction examples can produce a well-aligned model, emphasizing that data quality is more critical than quantity for efficient instruction tuning. Healthcare LLMs are susceptible to hallucinations (Longpre et al., 2023), highlighting the need of high data quality for evaluation—particularly in terms of factuality, safety, and diversity—prior to model tuning (Liu et al., 2023d). Leveraging SLMs for efficient and targeted data selection offers a promising strategy to address hallucination (Pezoulas et al., 2024; Shi et al., 2024) (see Fig. 5(a)). However, despite the use of SLM for SHDG (Liu et al., 2025a), the potential of SLMs in curating quality data remains largely unexplored. On device personalization. Patient-generated data often includes sensitive content, making cloud-based annotations undesirable or restricted in clinical settings. Local annotation is possible but must remain sparse to avoid user disruption, and limited device storage hampers large-scale fine-tuning. To this end, Qin et al. (2024a) proposed the on-device LLM personalization framework that selects and stores representative user data in a self-supervised manner, using minimal memory and occasional annotations (see Fig. 5(b)). It forms training mini-batches from streaming interactions without retaining the entire data stream. Experiments show superior personalization accuracy and fine-tuning speed over standard baselines. Ensembling SLMs. Cho et al. (2025) introduced a task-level ensemble technique that balances performance and costs by combining SLMs and LLMs, offering Pareto-optimal solutions. The power of ensembling SLMs for healthcare tasks is demonstrated by using multiple fine-tuned SLMs on different clinical report sections to ensure comprehensive coverage (Gondara et al., 2025) (see Fig. 5(c)). Combined with a LLM for conflict resolution, this approach achieves high accuracy in tumor classification from pathology reports. Distillation. Lightweight language models are vital for expanding mental healthcare access, especially in low-resource settings (Shi et al., 2024). With the use of knowledge distillation, high-performing cloud-based models can transfer knowledge to smaller edge models, boosting their efficiency and accuracy (see Fig. 5(d)). This enables real-time, personalized support on mobile devices while maintaining low resource demands, addressing global mental health gaps.
5.3. Assessment methods
Challenging the assumption that smaller models inherently compromise accuracy, recent research demonstrates that SLMs have the potential to perform effectively in specialized healthcare tasks (Zhang et al., 2025b). However, ethical concerns remain critical barriers, underscoring the need for rigorous validation across diverse patient populations to ensure responsible healthcare adoption. As ethical concerns grow, developers and clinicians must proactively mitigate risks to ensure trustworthiness in patient-centered use—especially as SLMs become increasingly personalized and mobile. Evaluating performance is crucial due to the sensitive nature of medical information, where inaccurate or unsafe responses could potentially cause harm. Healthcare-specific SLM evaluations prioritize domain knowledge correctness through accuracy and judgment, safety through hallucination control, and alignment with medical standards through trustworthiness, and better resource utilization through memory and energy efficiency.
Accuracy and Judgment: Chen et al. (2023) develop a novel, physician-curated adversarial question dataset rooted in real-world clinical scenarios, along with a comprehensive 17-metric evaluation framework designed to assess alignment and contextual relevance to clinical practice. Selene Mini is an open-source SLM designed for high-accuracy evaluation across diverse tasks [^14] https://huggingface.co/blog/AtlaAI/selene-1-mini. Trained with a principled blend of supervised and preference optimization on curated data, it outperforms 11 benchmarks of SLMs as a judge, demonstrating robust, domain-adaptable, and format-resistant performance for real-world deployment.
Table 7. Carbon emissions during the pretraining phase of various medical/healthcare LLMs, and the estimated number of trees required to offset one year of those emissions (assuming $0.021\,\text{tCO}_{2}$ absorbed per tree per year (Rainforest Trust, 2025)). Unmarked emission values are reported by legitimate sources, while values marked with $\approx$ are estimates obtained from base models.
[…]
AlpaCare (Zhang et al., 2023b) LLAMA-2 7B $\approx$0.01 1
Apollo (Wang et al., 2024a) Qwen-7B (v2) $\approx$30.0 1,429
BioGPT (Luo et al., 2022) GPT-2 Medium (345M) $\approx$0.10 5
BioMedLM (Bolton et al., 2024) GPT-2 (2.7B params) 2.00 (Bolton et al., 2024) 96
BioMistral (Labrak et al., 2024) Mistral-7B 0.074 (Labrak et al., 2024) 4
BioMistral-NLU (Fu et al., 2024) BioMistral-7B (fine-tuned) $\approx$0.02 1
CancerGPT (Li et al., 2024b) GPT 2 (124M) $\approx$0.001 1
CancerLLM (Li et al., 2024a) Mistral-7B $\approx$1.00 48
ChatCounsellor (Liu et al., 2023a) Vicuna-v1.3 7B (fine-tuned) $\approx$0.01 1
ChatDoctor (Li et al., 2023a) LLaMA-13B (fine-tuned) $\approx$0.01 1
ClinicalMamba (Yang et al., 2024a) Mamba (2.8B) $\approx$0.04 2
Med-$R^{2}$ (Lu et al., 2025) Qwen 2.5 (7B) $\approx$0.01 1
MedAlpaca (Han et al., 2023) LLAMA-2 7B $\approx$0.01 1
Meditron (Chen et al., 2023) LLaMA-2 (7B) (continued pretrain) 0.20151515https://huggingface.co/epfl-llm/meditron-7b 95
MentaLlama (Yang et al., 2024b) LLaMA-2 (7B/13B) $\approx$0.10 5
MentalQLM (Shi et al., 2024) Qwen1.5 $\approx$0.50 24
mhGPT (Kim et al., 2024a) GPT-based (1.98B) $\approx$0.10 5
PMC-LLAMA (Wu et al., [n. d.]) LLaMA-7B $\approx$0.08 4
RadPhi-2 (Ranjit et al., 2024a) Phi-2 (2.7B) 0.01 (Ranjit et al., 2024a) 1
RadPhi-3 (Ranjit et al., 2024b) Phi-3 (3.8B) $\approx$0.02 $\sim$1
Safety: A modular safety framework leveraging SLMs was proposed to detect and respond to harmful user queries, mitigating the computational overhead and performance compromises associated with embedding safety mechanisms directly within LLMs (Kwon et al., 2024). Employing SLMs as standalone safety modules in healthcare offers a scalable, efficient, and language-flexible approach to AI safety (Bhimani et al., 2025). Safety infuse medical validation among virtual health assistants to reduce their manipulation leveraged on controlled conversations through question generation, thereby deriving positivity in communications. A pervasive problem with increased reliability on language models is hallucinations which can be classified into factuality and faithfulness. Med-Pal’s robust safety alignment showcase its consistent refusal to engage with diverse adversarial prompts, including prompt injection, jailbreaking, misinformation, and harmful outputs—demonstrating effective defense against misuse in clinical dialogue systems (Elangovan et al., 2024). LookAhead Tuning enhances the safety of fine-tuned language models by incorporating partial answer previews during training, effectively preserving safety alignment while maintaining strong task performance with minimal computational overhead (Liu et al., 2025c). To address the critical gap in evaluating medical safety of LLMs, MedSafetyBench introduces the first benchmark grounded in medical ethics, revealing current models fall short and demonstrating that fine-tuning with this dataset enhances safety without compromising performance (Han et al., 2024).
Trustworthiness: Benchmarking LLMs in healthcare requires standardized, domain-specific metrics to ensure safety, accuracy, and ethical integrity while addressing data privacy, bias, and explainability (Budler et al., 2025; Aljohani et al., 2025). Truthfulness is critical in language models to prevent misinformation. Using the TruthfulQA benchmark, BioMistral-7B shows strong performance, outperforming GPT-3.5 Turbo by 4%, though no model excels across all categories, highlighting task-specific strengths and limitations (Labrak et al., 2024). Faithfulness in AI-generated medical information remains a critical concern, and recent efforts focus on understanding its causes, developing evaluation metrics, and proposing mitigation strategies to ensure factual accuracy across generative medical AI applications (Xie et al., 2023b).
Memory and energy efficiency: Energy efficiency is a high priority for SLMs, especially when used on battery-powered devices. MELODI is a novel framework and dataset designed to monitor and analyze energy consumption during LLM inference, revealing significant efficiency disparities and offering a foundation for sustainable, energy-conscious AI deployment (Husom et al., 2024). Response length significantly impact power consumption, and concise outputs can reduce energy usage. We measure the reduced environmental impact of healthcare SLM development through carbon footprints in Table 5.3. Greenhouse equivalences can be calculated with a well-establish approach [^16] https://www.epa.gov/energy/greenhouse-gas-equivalencies-calculator.
6. Discussion and conclusion
SLMs are emerging as valuable tools for assisting physicians and patients by enabling personalized treatment plans, accelerating diagnoses, and streamlining medical record analysis. Our investigation of existing NLP-driven SLMs in healthcare—focused on English-speaking countries—aims to support clinical decision-making across the care continuum. We report the search terms and inclusion-exclusion criteria for identifying relevant studies. The timeline of SLM development reveals advances in NLP tasks for specific clinical conditions, healthcare NLP tasks, medical NLP tasks, and doctor-patient conversations/mental health applications. We propose the taxonomies to investigate customization and tuning; the compression techniques for development of specialized SLMs; and outlook for healthcare professionals. We justify the need of SLM development in a high-stake application - healthcare informatics and curate the open-source datasets and models to facilitate future developments. Domain-specific tokenization has advanced healthcare language models by leveraging tools like UMLS, abbreviation disambiguation, and custom BPE methods to preserve medical terminology structure, improving classification accuracy and reducing training costs (Hosseini et al., 2024). The observation - input and output vocabularies exhibit distinct scaling behaviors in small models, raises open questions about optimal vocabulary construction, particularly in healthcare (Huang et al., 2025). How far such optimization can be achieved through data selection and curriculum learning remains an important area for future exploration. Concept-based reasoning in large concept models offers promising potential for future clinical concept extraction by enabling deeper understanding of full clinical statements, including negations and causal relationships (Barrault et al., 2024). We discovered the pressing need to benchmark data selection methods for constructing high quality datasets.
We discuss architectural foundations for model development and key advancements in SLMs. Given the central role of self-attention in lightweight model architecture, we present a comparative analysis of attention mechanisms based on time complexity to guide the efficient selection and deployment of models for healthcare NLP tasks. Although low-rank adaptation techniques have been explored for SLM development in the past, their application in clinical concept extraction—remains underexplored, presenting opportunities to enhance efficiency and adaptability in task-specific fine-tuning. Advancing trustworthy healthcare SLMs requires integrating human-in-the-loop reasoning, CoT, RAG, knowledge graphs, and emerging meta-cognitive frameworks to improve factual accuracy, contextual understanding, and intelligent decision-making across clinical domains (Sethi et al., 2025; Amugongo et al., 2024). With evolving CoT and RAG frameworks, there is an urgent need to investigate and advance in-context learning and reasoning mechanisms for knowledge-driven healthcare SLMs. Literature vouch for long training for small models (Bolton et al., 2024).
Noticably, many task-specific models were compressed using knowledge distillation, a few healthcare SLM models employ quantization, while pruning remains largely unexplored in this domain. SLMs play a pivotal role in addressing the limitations of LLMs in healthcare by offering efficient, scalable, and privacy-preserving solutions across various stages of clinical AI pipelines. From high-quality data curation and on-device personalization to ensembling and knowledge distillation, SLMs complement LLMs by enabling low-latency, cost-effective, and context-aware decision support.
Our goal is to provide healthcare professionals with practical insights on deploying SLMs in real-time applications, rather than explaining their internal workings. SLMs assist healthcare professionals by extracting key clinical concepts—such as social determinants of health, quality of life, and therapeutic interventions—from unstructured text. They support clinical decision-making through automated diagnosis assistance, personalized treatment recommendations, and summarization of lengthy patient records. Additionally, SLMs enhance real-time clinical decision support by identifying rare diseases, drug interactions, and aligning patient data with medical literature. SLMs support clinical research by structuring EHR data and monitoring adverse events, thereby reducing manual chart reviews and enhancing data usability. They automate clinical coding and billing by mapping free-text entries to standardized codes, minimizing administrative burden and claim errors. SLMs also enable clinical QA by extracting relevant insights from research literature and patient records, accelerating evidence-based investigations. SLMs have the potential to enhance patient-centered care by analyzing sentiment and monitoring mental health to detect risks like worsening depression. They empower virtual health assistants and triage systems, facilitating personalized engagement and real-time conversational support. Additionally, SLMs improve healthcare accessibility through Plain Language Adaptation (PLA) - simplifying complex medical information for individuals with limited health literacy.
We report our study on the rise of SLMs in healthcare at Github repository. We further investigate and report the reduction in carbon-emission and rise in sustainability due to evolving SLM in healthcare. Further in this section, our concluding remarks covers challenges and future research directions (§6.1), and the impact of healthcare SLMs (§6.2).
6.1. Challenges and research directions
The contrast between the theoretical capabilities and real world limitations is a fascinating space to explore. SLMs in NLP-driven healthcare informatics are facing a significant benchmarking challenge. Unlike general-domain NLP, healthcare informatics relies heavily on proprietary, heterogeneous data, limiting reproducibility and transparency. Public datasets offer partial solutions but often lack representativeness. As models grow in complexity and autonomy, evaluation frameworks must evolve to incorporate safety, ethics, and trustworthiness. Standardized benchmarks are essential to ensure consistent assessment, with key criteria including “doctor-like" communication—demonstrating medical expertise, sound reasoning, and clinically relevant diagnostic insight. Progress requires collaborative efforts, sustainable funding, regulatory alignment, and a workforce skilled in both informatics and health sciences. Edge deployment is critical for expanding access to AI tools in underserved regions, enabling point-of-care decision support, and reducing reliance on cloud-based APIs that may raise data privacy concerns. Interoperability is the model’s ability to seamlessly integrate and operate within the existing Health Information Exchanges (HIEs) from different sites. Another significant advantage of SLMs is their potential when used in multi-agent AI systems—collections of specialized agents that collaborate to solve complex tasks. Instead of relying on a single monolithic model like an LLM, multi-agent systems distribute tasks among several domain-specific SLMs. Future research must focus on developing modular interfaces, shared embedding spaces, and universal chaining frameworks that enable plug-and-play interoperability across models of varying sizes and capabilities. We posit the SLM development as a multi-dimensional solution for advancing NLP-centered healthcare informatics. Several active research directions in the development of generic SLMs are being effectively adapted to advance healthcare-specific SLM development. We enlist significant developments in healthcare as follows.
Human-in-the-loop systems integrate SLMs, LLMs, and expert oversight to ensure safe and effective deployment. SLMs assist with triage and uncertainty flagging, while LLMs manage complex synthesis. Embedding human feedback enhances trust and supports continual model refinement. Advancing flexible architectures, evaluation frameworks, and regulatory standards is essential, with future research focusing on efficient evaluators for clinical tasks, emphasizing factuality, safety, and uncertainty (Han et al., 2024; Xie et al., 2023b; Budler et al., 2025). Reducing manual chart review. As LLMs are preferentially aligned with human-preferred responses and surpass human capabilities, the feasibility and reliability of human evaluation diminish, introducing a weak-to-strong generalization, wherein smaller, less capable models supervise and guide the alignment of more powerful models (Burns et al., 2023). This paradigm has potential to reduce manual chart review efforts, enabling scalable, cost-effective, safe and factual preference alignment that exceed clinician expertise. Summarizing medical reports. SLMs perform high-quality abstractive summarization by distilling structured rationales from LLMs (Wang et al., 2025a). These rationales encapsulate important structural and semantic cues from the document, broken into high-level concepts, text-spans that support those themes, and relationships among key entities in the document. These structured rationales serve as interpretable reasoning traces that guide summary generation, similar to CoT prompting but in a structured, modular format. SLMs are rationale-aware, lightweight summarizers that are trained using intermediate reasoning signals setting a new paradigm for transparent and cost-effective summarization. Bias detection and de-biasing models. Because SLMs are often fine-tuned on relatively small clinical corpora (e.g., institutional EHRs or specialty-specific datasets), they are at high risk of inheriting data biases or overfitting to local linguistic patterns (Sapkota et al., 2025). This can manifest in poor generalization to diverse patient populations, inaccurate risk stratification, or propagation of health disparities in AI-driven decision-making. Future research must incorporate domain-aware debiasing techniques such as adversarial training, fairness-aware objective functions, and counterfactual data augmentation. Rigorous bias audits and alignment assessments should become standard in the development lifecycle of healthcare SLMs. Automated Machine Learning (AutoML). While existing model compression techniques have shown significant progress, they continue to rely heavily on manual design—particularly in defining suitable student architectures for knowledge distillation, which demands considerable human effort. The application of AutoML to LLM remains limited due to the significant computational burden involved in scaling down models with billions of parameters. SLMs address this challenge by evaluating architectural components like model depth and attention head width within a reduced parameter space (Liu et al., 2024e). Shen et al. (2024) reduce the search complexity by incorporating optimized initialization strategies, accelerating convergence. Therefore, the models must integrate AutoML approaches—meta-learning and neural architecture search, that offer a promising direction to mitigate this dependency. This will enhance compression efficiency without compromising performance.
6.2. Impact of SLMs
SLMs offer infrastructure-friendly deployment on edge devices using optimized hardware such as NVIDIA Jetson or Qualcomm AI accelerators. Software optimization techniques (distillation, quantization and pruning) significantly lower computational demands, enabling model adoption without costly hardware investments. However, challenges such as frequent model updates and deployment across heterogeneous devices may introduce new scalability and maintenance complexities. Given these considerations, investigating the impact of model development is necessary to support responsible application. Environmental and economic trade-offs: Training GPT-3 with 175 billion parameters consumed 1287 MWh of electricity, and resulted in carbon emissions of 502 metric tons of carbon, equivalent to driving 112 gasoline powered cars for a year (Patterson et al., 2021). The environmental impact of AI is an escalating concern (Rillig et al., 2023) where SLMs offer a partial solution by reducing the training and inference energy consumption by up to 80% (Liu and Yin, 2024). This reduction directly translates into a smaller carbon footprint, making SLMs a more sustainable choice as evident from Table 5.3. This efficiency is valuable in healthcare, aligning with the growing movement toward green AI, where the need of real-time solutions must be balanced with sustainability goals. The future directions for environmentally conscious green (small/large) language model development in healthcare promotes compression techniques discussed in §4. Ethical and Social Impact. The integration of SLMs into clinical workflows alleviates provider burden while promoting equitable access, reducing digital disparities, and enhancing patient-centered care. Their lightweight nature enables deployment in low-resource settings, including rural clinics, mobile health units, and edge devices, thereby extending AI-driven support to underserved populations. By lowering the technical and financial barriers associated with large models, SLMs democratize the use of advanced NLP tools across diverse healthcare institutions (Mittal et al., 2024). Additionally, SLMs support data de-identification by enabling efficient, on-device removal of personally identifiable information (PII) from clinical texts, facilitating privacy-preserving data sharing and compliance with healthcare regulations. SLMs can support patient friendly adoption through plain language adaptation, and empower patients with timely, comprehensible information. However, addressing fairness, bias, and equitable access remains critical (Sapkota et al., 2025). While the narrower focus of model training datasets allows for more effective bias filtering, it also risks excluding underrepresented groups or contexts. Therefore, careful dataset design is essential to ensure inclusive and responsible AI deployment. Synthesis and Perspectives. The goal of SLMs is to prioritize efficiency and accessibility. While, they mitigate environmental, economic, and ethical challenges associated with LLMs, the success of these methods largely depend on interdisciplinary collaborations, for example, hardware engineers optimizing for energy efficiency, clinicians validating medical tools, and policymakers ensuring equitable access. Future work in this direction requires standardized benchmarks for fairness and environmental impact, a step toward sustainable, human-centric AI.
6.3. Concluding remarks
This work underscores the transformative potential of SLMs in democratizing AI for healthcare, advocating for models that harmonize performance and ethical responsibility with limited resources. Future efforts must prioritize clinician-AI collaboration and regulatory alignment to unlock scalable, human-centric solutions for NLP-centered clinical applications in English speaking territories. Although LLMs possess extensive medical knowledge, smaller, specialized clinical LLMs outperform them, even when fine-tuned on limited annotated data (Zhou et al., 2023b). However, generic SLM have potential to achieve comparable performance to LLMs in targeted use cases while reducing computational costs, emphasizing the sustainable model development (Allal et al., 2025). However, whether task-specific SLM in healthcare achieve comparable performance to LLMs or not, is still an open research question (Labrak et al., 2024; Fu et al., 2024; Zhang et al., 2025b). Efforts have been made to explore medical QA, language understanding and mental health analysis; a huge gap lies in model development for clinical concept extraction and report summarization. Current research on SLMs remains nascent, with limited exploration of longitudinal real-world deployments.
While this study focuses on NLP-centered healthcare informatics for English speaking territories, we acknowledge the growing importance of multilingual and non-English models, which is beyond our current scope. Our review excludes multimodal research and the model lacking peer-reviewed evidence. Moving forward, we aim to maintain an up-to-date repository on GitHub to reflect ongoing developments in this evolving field.
Acknowledgements
This study was supported by NIH (National Institutes of Health) R01 AG068007.